OpenStudy (anonymous):

What is the solution to the equation? 4(3 + 7) = 4 • x + 4 • 7 A. x = 3 B. x = 4 C. x = 7 D. x = 10

OpenStudy (liv1234):

PEMDAS

OpenStudy (landon34):

ill solve it.

OpenStudy (liv1234):

Parentheses Exponents Multiplication Division Addition Subtraction

OpenStudy (landon34):

hold on, ill explain it to you in steps.

OpenStudy (maurticewilliams):

i was gonna help but it looks like someone else already got to it

OpenStudy (liv1234):

(3+7)=10*4=40

OpenStudy (thecuteone25):

the answer is X=3

OpenStudy (liv1234):

No direct answers.^

OpenStudy (landon34):

Simplifying 4(3 + 7) = 4x + 4 * 7 Combine like terms: 3 + 7 = 10 4(10) = 4x + 4 * 7 Multiply 4 * 10 40 = 4x + 4 * 7 Multiply 4 * 7 40 = 4x + 28 Reorder the terms: 40 = 28 + 4x Solving 40 = 28 + 4x Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-4x' to each side of the equation. 40 + -4x = 28 + 4x + -4x Combine like terms: 4x + -4x = 0 40 + -4x = 28 + 0 40 + -4x = 28 Add '-40' to each side of the equation. 40 + -40 + -4x = 28 + -40 Combine like terms: 40 + -40 = 0 0 + -4x = 28 + -40 -4x = 28 + -40 Combine like terms: 28 + -40 = -12 -4x = -12 Divide each side by '-4'. x = 3 Simplifying x = 3

OpenStudy (f_jayyy):

The answer is A

OpenStudy (landon34):

I hope that helps, your welcome.

OpenStudy (anonymous):

Thanks people!

OpenStudy (landon34):

did you see my explanation?

OpenStudy (anonymous):

Yes thankyou i understand :) adios

OpenStudy (landon34):

ok.